An analyst computes the beta of the computer company WinDoze as 1.7 and the standard error of the estimate as 0.3. What is the 95 percent confidence interval for the calculated beta? how your work with explanation!
Answer #1

margin of error = z-score * standard error

= 1.96 * 0.3

= 0.588

lower bound = mean - margin of error

= 1.7 - 0.588

= 1.112

Upper bound = mean + margin of error

= 1.7 + 0.588

= 2.288

hence the interval is [1.112,2.288]